OpenStudy (anonymous):

The lengths of the sides of a triangle form three consecutive odd numbers. The perimeter is 21 centimeters. What is the length of the longest side?

OpenStudy (anonymous):

Consecutive odd numbers differ by 2. n + (n + 2) + (n + 4) = 21 n + n + 2 + n + 4 = 21 3n+6=21 3n=21-6=15 n=5 5, 5+2=7, 5+4=9 5+7+9=21

OpenStudy (anonymous):

Small thing (very small). It's a slicker operation if you call them (n-2) , n, (n+2) I mean, obviously it makes no difference, but I prefer it :D

OpenStudy (anonymous):

Yes, the twos cancel. ------- Did you review my solution to your "fun" problem the other day?

OpenStudy (anonymous):

no, sorry :( I'll try and find that thread again

OpenStudy (anonymous):

A real shame :( You (I think) misread the fractions 1/3 1/5 1/7 etc as 13, 15, 17. I think that is the cause of the difference. I assume yours are correct to the (slightly different) problem though.

OpenStudy (anonymous):

I will redo it with your fractions. Thanks for responding.

OpenStudy (anonymous):

I look forward to it

OpenStudy (anonymous):

After my last post above, I remember taking a screen capture of the problems that you posted on this math site so that I would not make any errors in the equations to be solved. Refer to the attachment, a screen capture. I will check the equations carefully again.

OpenStudy (anonymous):

Oh, weird, your computer doesn't render LaTeX on this site :/

OpenStudy (anonymous):

What do you see below? \[\frac{4}{5} \] ? Because that renders as a fraction (4/5), do you see 45?

OpenStudy (anonymous):

I see 45 on the screen. Unable today to enter x^2 and have complained about it. x^2 is present as 2x in their window. I'll try another computer or use a virtuaized version of Ubuntu Linux.

OpenStudy (anonymous):

:( Sorry for any confusion. For the record the full question is in this image: one jpeg and one tif